OpenStudy (anonymous):

If a .10 M solution of a colored substance has a maximum absorbance at 500 nm and an absorbance of .26 at this wavelength, what will be the measured absorbance of a .20 M solution at 500 nm?

OpenStudy (aaronq):

\(Abs \propto concentration\) using the beer-lambert law: \(\dfrac{Abs_1}{C_1}=\dfrac{Abs_2}{C_2}\)

OpenStudy (anonymous):

I have not learned that formula yet

OpenStudy (aaronq):

the actual formula is: \(Abs=\epsilon*l*C\), i just simplified it

OpenStudy (anonymous):

I do not know that formula unless it is like -log(1-decimL)

OpenStudy (aaronq):

it's just like Charles' Law: \(\dfrac{V_1}{T_1}=\dfrac{V_2}{T_2}\)

OpenStudy (anonymous):

this question is just to hard I ended up with a number that is totally wrong it was like -12.3784908

OpenStudy (aaronq):

plug your numbers into the formula (in my first post).

OpenStudy (anonymous):

is there a fractction

OpenStudy (anonymous):

500/.26

OpenStudy (anonymous):

do I divide it by 50/.2

OpenStudy (anonymous):



OpenStudy (anonymous):

i don't know where the numbers go?

OpenStudy (aaronq):

Abs = absorbance C= concentration you don't need to use the wavelength (500 nm)

OpenStudy (anonymous):

so you are saying  equals